OpenStudy (anonymous):

Solve for x. 8 – 3x – x = 12

OpenStudy (anonymous):

*ahem** cup?

OpenStudy (anonymous):

why

OpenStudy (anonymous):

i love these

OpenStudy (anonymous):

Sorry, I want to play chess with someone else, I need to go pet sit my brothers dog

OpenStudy (anonymous):

8 - 12 = 4x x = -4/4 = -1 i think

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

Simplifying 8 + -3x + -1x = 12 Combine like terms: -3x + -1x = -4x 8 + -4x = 12 Solving 8 + -4x = 12 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-8' to each side of the equation. 8 + -8 + -4x = 12 + -8 Combine like terms: 8 + -8 = 0 0 + -4x = 12 + -8 -4x = 12 + -8 Combine like terms: 12 + -8 = 4 -4x = 4 Divide each side by '-4'. x = -1 Simplifying x = -1 yep thats right

OpenStudy (anonymous):

anymore and welcome